# Foundations of ecological and evolutionary change

### v1

@@ Abstract
Biological evolution is realised through the same mechanisms of birth and death that underlie change in population density. The deep interdependence between ecology and evolution is well-established, but much theory in each discipline has been developed in isolation. While recent work has accomplished eco-evolutionary synthesis, a gap remains between the logical foundations of ecology and evolution. We bridge this gap with a new equation that defines a summed value for a characteristic across individuals in a population, from which the fundamental equations of population ecology and evolutionary biology (the Price equation) are derived. We thereby unify the fundamental equations of population ecology and biological evolution under a general framework. Our unification further demonstrates the equivalence between mean population growth rate and evolutionary fitness, shows how ecological and evolutionary change are reflected in the first and second statistical moments of fitness, respectively, and links this change to ecosystem function.
@@ Introduction
We therefore begin with the most fundamental axioms underlying the ecology and evolution of living systems (Rice 2004; Rice and Papadopoulos 2009). In such systems, diversity is discontinuous and can be defined in terms of discrete entities (Dobzhansky 1970). Our framework is general enough that entities can be anything discrete, but we will focus on each entity $i$ as an individual organism. Individuals give rise to new individuals through birth such that $\beta_{i}$ is the number of births attributable to $i$. Individuals are removed from the population through death such that $\delta_{i}$ is an indicator variable that takes a value of 1 (death of $i$) or 0 (persistence of $i$). All individuals are defined by some characteristic $z_{i}$, and $\Delta z_{i}$ defines any change in $z_{i}$ from one time step $t$ to the next $t+1$. The total number of individuals in the population at $t$ is $N$. From this foundation, we can define $\Omega$ to be a summed characteristic across $N$ entities,
@@
We can now interpret $\Omega$ as the population size in $t+1$, $N_{t+1}$ and note that $N\beta$, $N\delta$, and $N$ are the total births, total deaths, and size of the population at $t$, respectively,
@@
We can break this down further and multiply each side by $1/N$,
@@
Now we must consider the total conservation of probability (Frank 2015, 2016). In eqn 7, $\Omega$ is the total sum trait values ($z_{i}$) across the entire population at $t+1$ divided by the number of individuals ($N$) in the population at $t$. But the size of the population can change from $t$ to $t+1$. To recover mean trait change for the Price equation (and therefore conserve total probability), we need to account for this change in population size. The left-hand side of eqn 7 describes contributions to the sum trait value from the new population at $t+1$. But we cannot treat $\Omega/N$ as the mean of $z$ at $t+1$ ($\bar{z}^{\prime}$) because we need to weigh $N$ by the mean fitness of the population at $t$ to account for any change in population size from $t$ to $t+1$. We need the mean $z$ across $N$ at $t$.
For example, if mean fitness at $t$ is 2 (i.e., $\bar{w}=2$), then half as many individuals will have contributed to $\Omega$ in $t+1$ than would have if $\bar{w}=1$ (i.e., there are $N$ individuals at $t$ and $2N$ individuals at $t+1$). We therefore need to multiply the mean trait value of $\bar{z}^{\prime}$ (at $t+1$) by the mean fitness of $\bar{w}$ at $t$ to recover the mean contribution of the $N$ individuals at $t$ to the total $\Omega$. Consequently,
@@
Equation 8 conserves the total probability and now clarifies $\Omega$ as a summed trait value, which equals expected population growth at $t$ (note $\lambda=\bar{w}$) times mean trait value at $t+1$. We can therefore rewrite,
@@ Discussion
Conceptual unification has a critical role in advancing scientific theory. Showing how apparently disparate phenomena follow from the same shared principles reveals what is most fundamental about the natural world and provides a foundation for further scientific investigation and the construction of more coherent and predictive models. Our eqn 1 provides a foundation for the conceptual unification of the fundamental equations of population ecology and evolution.
@@
The Price equation provides a complete and exact description of evolution in any closed evolving system (Price 1970; Frank 2012). It is derived by rearranging the mathematical notation defining changes in the frequencies and characteristics of any type of entity (e.g., individuals, alleles, Price 1970; Gardner 2008; Luque 2017). This derivation partitions total characteristic change into different components, making it possible to isolate evolutionary mechanisms (e.g., selection) and levels of biological organisation (e.g., group, individual, Frank 1995, 2012; Kerr and Godfrey-Smith 2009; Luque 2017; Okasha and Otsuka 2020). Because of its abstract nature and lack of any system specific assumptions, the Price equation makes no predictions about what will happen in any particular system (Gardner 2020). Its role is not to predict, but to formally and completely define and separate components of evolutionary change. The same is true of the general equation for population change (eqns 2 and 3), at least as we have used it here where it serves to define what population growth means in ecology. This equation formally and completely describes population change in terms of births and deaths. In eqn 1, we therefore have a fundamental equation from which we can derive complete ecological and evolutionary change in any closed biological population.
Our unification and derivation of fundamental equations of ecology and evolution recovers and formalises well-known properties of biological systems. The expressions $\beta_{i}-\delta_{i}+1$ and $z_{i}-\Delta z_{i}$ can be interpreted to represent the ecological and evolutionary components of a system, respectively. When focused only on change in population size, we set $z_{i}=1$ and $\Delta z_{i}=0$, then we can derive the fundamentals of population ecology without considering evolution. Interestingly, the same is not true when focused only on evolutionary change. If we set $\beta_{i}=0$ and $\delta_{i}=0$ (no birth or death), then there is no variance in fitness ($w_{i}=1$ for all individuals), and evolution does not occur (note that evolution can still occur in a population at equilibrium where the birth rate and death rate are the same). Fundamentally, evolutionary change is therefore enacted through ecological change, i.e., individual births and deaths, but ecological change does not necessarily imply evolution. And while the link to population change is intuitive from the structure of eqn 1, the link to the Price equation is much less intuitive, especially after eqn 5 in our derivation. The Price equation describes change in any evolving system, not just biological systems (Fox 2006; Brantingham 2007; Bell, Richerson, and McElreath 2009; MacCallum et al. 2012; Frank 2015). It is perhaps unsurprising that a derivation of ecological and evolutionary change from a common starting point requires relatively more labour to recover the Price equation.
@@
Our unification demonstrates the equivalence between the finite rate of increase $\lambda$ (Box 1, Gotelli 2001) and population mean evolutionary fitness $\bar{w}$ (Box 2). The population growth equation $N_{t+1}=N_{t}\lambda$ could therefore be rewritten as $N_{t+1}=N_{t}\bar{w}$. We do not claim to be the first to notice this equivalence between population growth and mean population fitness. Indeed, the relationship between population growth rate and evolutionary fitness has been proposed and applied many times before (e.g., Fisher 1930; Charlesworth 1980; Lande 1982; Roff 2008; Lion 2018). For example, Lande (1982) explicitly concludes mean absolute fitness per unit time is $\bar{w}=e^{r}$, where $r=ln(\lambda)$. We show this from first principles and clarify the relationship between fitness and population growth. Over an arbitrary length of time, fitness is properly defined as $w_{i}=\beta_{i}-\delta_{i}+1$. Over an individual’s lifetime (which, by definition, includes death), fitness is therefore $\beta_{i}$. Interestingly, the rate of change in ecology and evolution are reflected in the first and second statistical moments of fitness, respectively. Population growth rate reflects mean population fitness $\bar{w}$, while the rate of evolutionary change reflects the variance in fitness $\mathrm{Var}(w)$ (i.e., Fisher’s fundamental theorem, Frank 1997; Rice 2004; Queller 2017).
@@
All of these connections may also help to explain, at least partially, some of the success of classical population genetic models. For decades, population genetics (and to some extent quantitative genetics) has been accused of being a reductionistic view of evolution, reducing everything to changes in allele frequencies and abstracting away from individuals and their environments (e.g., the ecological interactions, MacColl 2011). This has been a line of argumentation by some defenders of the so-called Extended Synthesis (Pigliucci 2009), especially in relation to niche construction (Odling-Smee, Laland, and Feldman 2003). Famously, Mayr (1959) characterized population genetics as a simple input and output of genes, analogous to “the adding of certain beans to a beanbag and the withdrawing of others” (also called “beanbag genetics”).
@@
Historical critics of population genetics could not articulate a clear explanation for why it works so well despite all of its idealisations and simplifications. From the Price equation, we are able to recover classical population and quantitative genetic models (Queller 2017) and develop new ones (Rice 2004, 2020; Luque 2017; Lion 2018). Our eqn 1 contains ecology at its core, and we show how the Price equation logically follows from it. We therefore conclude that population and quantitative genetics equations contain ecology (no matter how hidden) as a mathematical necessity. And since we cannot have evolution without ecology, the ecological nature of evolution is implicit in population and quantitative genetic models. A classical sign of scientific progress is the ability to connect disparate theories and models to show how empirical and theoretical models are logical (mathematical) consequences of more fundamental ones (Nagel 1961).
@@
We have shown that we can derive the fundamental equations of population ecology and biological evolution from a single unifying equation. Lastly, we propose our eqn 1 as a potential starting point for defining ecosystem function and further conceptual unification between ecosystem function, ecology, and evolution. The Price equation has previously been used to investigate ecosystem function (Loreau and Hector 2001; Fox 2006), but not with any attempt towards conceptual unification with evolutionary biology. For example, Fox (2006) applied the abstract properties of the Price equation to partition total change in ecosystem function into separate components attributable to species richness, species composition, and context dependent effects. This approach provides a framework for comparing the effects of biodiversity on ecosystem function in empirical systems (e.g., Fox 2006; Winfree et al. 2015; Mateo-Tomás et al. 2017). Instead, our eqn 1 defines $\Omega$ as total ecosystem function contributed by a focal population (e.g., biomass production, decomposition, photosynthetic rate). It is therefore possible to investigate ecosystem function, population ecology, and population evolution from the same shared framework.

### v2

@@ Abstract
Biological evolution is realised through the same mechanisms of birth and death that underlie change in population density. The deep interdependence between ecology and evolution is well-established, but much theory in each discipline has been developed in isolation. While recent work has accomplished eco-evolutionary synthesis, a gap remains between the logical foundations of ecology and evolution. We bridge this gap with a new equation that defines a summed value for a characteristic across individuals in a population, from which the fundamental equations of population ecology and evolutionary biology (the Price equation) are derived. We thereby unify the fundamental equations of population ecology and biological evolution under a general framework. Our unification further demonstrates the equivalence between mean population growth rate and evolutionary fitness, shows how ecological and evolutionary change are reflected in the first and second statistical moments of fitness, respectively, and links this change to ecosystem function. Finally, we outline how our proposed framework can be used to unify social evolution and density-dependent population growth.
@@ Introduction
We therefore begin with the most fundamental axioms underlying the ecology and evolution of living systems (Rice 2004; Rice and Papadopoulos 2009). In such systems, diversity is discontinuous and can be defined in terms of discrete entities (Dobzhansky 1970). Our framework is general enough that entities can be anything discrete, but we will focus on each entity $i$ as an individual organism. Individuals give rise to new individuals through birth such that $\beta_{i}$ is the number of births attributable to $i$. Individuals are removed from the population through death such that $\delta_{i}$ is an indicator variable that takes a value of 1 (death of $i$) or 0 (persistence of $i$). All individuals are defined by some characteristic $z_{i}$, and $\Delta z_{i}$ defines any change in $z_{i}$ from one time step $t$ to the next $t+1$. The total number of individuals in the population is $N$. From this foundation, we can define $\Omega$ to be a summed characteristic across $N$ entities,
@@
We can now interpret $\Omega$ as the population size at $t+1$, $N_{t+1}$ and note that $N\beta$, $N\delta$, and $N$ are the total births, total deaths, and size of the population at $t$, respectively,
@@
We can break eqn 5 down further and multiply each side by $1/N$,
@@
Now we must consider the total conservation of probability (Frank 2015, 2016). In eqn 7, $\Omega$ is the total sum trait values ($z_{i}$) across the entire population at $t+1$ divided by the number of individuals ($N$) in the population at $t$. But the size of the population can change from $t$ to $t+1$. To recover mean trait change for the Price equation (and therefore conserve total probability), we need to account for this change in population size. The left-hand side of eqn 7 describes contributions to the sum trait value from the new population at $t+1$. But we cannot treat $\Omega/N$ as the mean of $z$ at $t+1$ ($\bar{z}^{\prime}$) because we need to weigh $N$ by the mean fitness of the population at $t$ to account for any change in population size from $t$ to $t+1$.
For example, if mean fitness at $t$ is 2 (i.e., $\bar{w}=2$), then half as many individuals will have contributed to $\Omega$ in $t+1$ than would have if $\bar{w}=1$ (i.e., there are $N$ individuals at $t$ and $2N$ individuals at $t+1$). We therefore need to multiply the mean trait value $\bar{z}^{\prime}$ (at $t+1$) by the mean fitness $\bar{w}$ (at $t$) to recover the mean contribution of the $N$ individuals at $t$ to the total $\Omega$. Consequently,
@@
Equation 8 conserves the total probability and now clarifies $\Omega$ as a summed trait value, which equals expected population growth at $t$ times mean trait value at $t+1$. This is consistent with the population ecology derivation from the previous section where $z_{i}=1$ by definition, and $\Omega=N_{t+1}$ (note $\lambda=\bar{w}$). We can therefore rewrite eqn 7,
@@ Discussion
A classical sign of scientific progress is the ability to connect disparate theories and models to show how empirical and theoretical models are logical (mathematical) consequences of more fundamental ones (Nagel 1961). Conceptual unification therefore has a critical role in advancing scientific theory (Morrison 2000). Showing how apparently disparate phenomena follow from the same shared principles reveals what is most fundamental about the natural world and provides a foundation for further scientific investigation and the construction of more coherent and predictive models. Our eqn 1 provides a foundation for the conceptual unification of the fundamental equations of population ecology and evolution.
@@
The Price equation provides a complete and exact description of evolution in any closed evolving system (Price 1970; Frank 2012). It is derived by rearranging the mathematical notation defining changes in the frequencies and characteristics of any type of entity (e.g., individuals, alleles, Price 1970; Gardner 2008; Luque 2017). This derivation partitions total characteristic change into different components, making it possible to isolate evolutionary mechanisms (e.g., selection) and levels of biological organisation (e.g., group, individual, Frank 1995, 2012; Kerr and Godfrey-Smith 2009; Luque 2017; Okasha and Otsuka 2020). Because of its abstract nature and lack of any system-specific assumptions, the Price equation makes no predictions about what will happen in any particular system (Gardner 2020). Its role is not to predict, but to formally and completely define and separate components of evolutionary change. The same is true of the general equation for population change (eqns 2 and 3), at least as we have used it here where it serves to define what population growth means in ecology. This equation formally and completely describes population change in terms of births and deaths. In eqn 1, we therefore have a fundamental equation from which we can derive complete ecological and evolutionary change in any closed biological population. We anticipate that this will be useful for eco-evolutionary theory in the same way that the Price equation is useful for evolutionary theory: facilitating specific model development and identifying new conceptual insights, unresolved errors, and sources of model disagreements.
@@
Our unification demonstrates the equivalence between the finite rate of increase $\lambda$ (Box 1, Gotelli 2001) and population mean evolutionary fitness $\bar{w}$ (Box 2). The population growth equation $N_{t+1}=N_{t}\lambda$ could therefore be rewritten as $N_{t+1}=N_{t}\bar{w}$. We do not claim to be the first to notice this equivalence between population growth and mean population fitness. Indeed, the relationship between population growth rate and evolutionary fitness has been proposed and applied many times before (e.g., Fisher 1930; Charlesworth 1980; Lande 1982; Roff 2008; Lion 2018). For example, Lande (1982) explicitly concludes mean absolute fitness per unit time is $\bar{w}=e^{r}$, where $r=ln(\lambda)$. We show this from first principles and clarify the relationship between fitness and population growth. Over an arbitrary length of time, fitness is properly defined as $w_{i}=\beta_{i}-\delta_{i}+1$. Over an individual’s lifetime (which, by definition, includes death), fitness is therefore $\beta_{i}$. Interestingly, the rate of change in ecology and evolution are reflected in the first and second statistical moments of fitness, respectively. Population growth rate reflects mean fitness $\bar{w}$, while the rate of evolutionary change reflects the variance in fitness $\mathrm{Var}(w)$ (i.e., Fisher’s fundamental theorem, Frank 1997; Rice 2004; Queller 2017).
@@
Our unification may also help explain, at least partially, some of the success of classical population genetic models. For decades, population genetics (and to some extent quantitative genetics) has been accused of being a reductionistic view of evolution, reducing everything to changes in allele frequencies and abstracting away from individuals and their environments (e.g., the ecological interactions, MacColl 2011). This has been a line of argumentation by some defenders of the so-called Extended Synthesis (Pigliucci 2009), especially in relation to niche construction (Odling-Smee, Laland, and Feldman 2003). Famously, Mayr (1959) characterized population genetics as a simple input and output of genes, analogous to “the adding of certain beans to a beanbag and the withdrawing of others” (also called “beanbag genetics”).
@@
Historical critics of population genetics could not articulate a clear explanation for why it works so well despite all of its idealisations and simplifications. From the Price equation, we are able to recover classical population and quantitative genetic models (Queller 2017) and develop new ones (Rice 2004, 2020; Luque 2017; Lion 2018). Our eqn 1 contains ecology at its core, and we show how the Price equation logically follows from it after accounting for absolute population growth (eqn 8). We therefore conclude that population and quantitative genetic equations contain ecology (no matter how hidden), and the ecological nature of evolution is implicit in population and quantitative genetic models.
@@
Our fundamental equation is complete and exact. It therefore implicitly includes any effects of density dependence on population growth (see Box 1), or any social effects on evolutionary change (see Box 2). We can make both of these effects explicit with the same partition in the summation on the right-hand side of eqn 1. To do this, we multiply the expression $\left(\beta_{i}-\delta_{i}+1\right)\left(z_{i}+\Delta z_{i}\right)$ by $\left(1-\sum_{j=1}^{N}a_{ij}\left(z_{j}+\Delta z_{j}\right)\right)$, where $j$ indexes the same individuals as $i$, and $a_{ij}$ defines the effect $j$ has on the contribution of $i$ to $\Omega$. In evolutionary terms, $a_{ij}$ defines how $j$ modulates the fitness of $i$ ($w_{i}$). In ecological terms, $a_{ij}$ defines how $j$ modulates the contribution of $i$ to total population growth. If we define $\alpha$ as the summed effect of all individuals on a focal $i$ (i.e., $\alpha=\sum_{j=1}^{N}a_{ij}$), then we can partition ecological change into density-independent and density-dependent effects and derive a discrete time logistic growth equation. We save these derivations for a future investigations focused on evolutionary and community ecology.
We have shown that we can derive the fundamental equations of population ecology and biological evolution from a single unifying equation. Lastly, we propose our eqn 1 as a potential starting point for defining ecosystem function and further conceptual unification between ecology, evolution, and ecosystem function. The Price equation has previously been used to investigate ecosystem function (Loreau and Hector 2001; Fox 2006), but not with any attempt towards conceptual unification with evolutionary biology. For example, Fox (2006) applied the abstract properties of the Price equation to partition total change in ecosystem function into separate components attributable to species richness, species composition, and context dependent effects. This approach provides a framework for comparing the effects of biodiversity on ecosystem function in empirical systems (e.g., Fox 2006; Winfree et al. 2015; Mateo-Tomás et al. 2017). Instead, our eqn 1 defines $\Omega$ as total ecosystem function contributed by a focal population (e.g., biomass production, decomposition, photosynthetic rate). It is therefore possible to investigate ecological, evolutionary, and ecosystem function change from the same shared framework.